# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ 7 Conclusion
In this paper, we model learning from imperfect human feedback as a corrupted continuous dueling bandit. We highlight the fundamental difficulty of LIHF by developing a regret lower bound. Then we show that by carefully tuning the learning rates of gradient-based methods, we can gracefully trade off learning efficiency for adversarial robustness. Our theoretical predictions also hold on real-world dataset. For future direction, it is interesting to generalize our setting to include contextual information and design algorithms robust to agnostic corruption. In addition, it is also an interesting open problem to design algorithms performs best-of-both-worlds, i.e. performing optimally both with and without agnostic corruption, in continuous action space.
A Connection between Imperfect User and Generalized Learnability
In this section, we establish the connection between adversarial corruption, corruption induced by generalized learnability, and corruption induced by $\rho$-imperfect user. Broadly speaking, holding total corruption budget $C$ constant, corruption induced by generalized learnability is a specific type of arbitrary corruption. Furthermore, if the proposed action pairs exhibit sufficient diversity, corruption induced by $\rho$-imperfect user is equivalent to corruption induced by generalized learnability.
In the following, we introduce the concept of the generalized learnability of the utility function $\mu$ in the scenario of pairwise comparison and the associated corruption induced throughout the user’s learning process.
Definition A.1 (Corruption induced by Generalized Learnability).
There exists an algorithm $\mathcal{G}$ such that, given any arbitrary $t$ pairs of actions $\{a_{s},a^{\prime}_{s}\}^{t}_{s=1}$ and the associated utility differences $\{y_{s}\}^{t}_{s=1}$, where $y_{s}:=\mu(a_{s})-\mu_{(}a^{\prime}_{s})$, $\mathcal{G}$ can output an estimated utility function $\mu_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ of the true utility function $\mu:\mathbb{R}^{d}\rightarrow\mathbb{R}$ such that for any action pair $(a,a^{\prime})$ the following equation holds
$|c_{t}(a,a^{\prime})|=|\mu_{t}(a)-\mu(a)-(\mu_{t}(a^{\prime})-\mu(a^{\prime}))|\leq C_{0}\left(\|a-a^{\prime}\|^{2}_{\bar{V}_{t}^{-1}}\right)^{1-\rho}.$
$C_{0}$ is some positive constant, $\rho\in[0,1]$, and $\bar{V}_{t}:=\lambda\mathbb{I}_{d}+\sum^{t}_{s=1}(a_{s}-a^{\prime}_{s})(a_{s}-a^{\prime}_{s})^{\top}$, where $\lambda$ is some positive constant to ensure that $\bar{V}_{t}$ is full rank.
Def. A.1 is a variant of the generalized learnability assumption proposed by Wang et al. (2023) and Yao et al. (2022a). Essentially, the parameter $\rho$ can be viewed as the user’s irrationality level. A higher value of $\rho$ implies lower learning speed of the user’s utility function $\mu$, hence a larger magnitude of the total induced corruption. Let’s consider arbitrary corruption with total corruption budget $C=\Theta(T^{\rho})$. Using Lemma A.1, we can show that $c_{t}(a,a^{\prime})$ induced by the generalized learnability of $\mu$ satisfies $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq O(T^{\rho})$, implies that it is a specific type of arbitrary corruption.
In the scenario when $\lambda_{\min}(\bar{V}_{t})$ increases in the order of $\Theta(t)$, $(\|a-a^{\prime}\|^{2}_{\bar{V}^{-1}_{t}})^{1-\rho}$ has the order $\Theta(t^{\rho-1})$, which implies there exists a positive constant $C_{\kappa}$ that at round $t$, given arbitrarily proposed action pairs $(a,a^{\prime})$, the magnitude of corruption $|c_{t}(a,a^{\prime})|\leq C_{\kappa}t^{\rho-1}$, which means that corruption induced by a $\rho$-imperfect user coincides with corruption induced by the generalized learnability of $\mu$.
We would like to highlight that the condition which $\lambda_{\min}(\bar{V}_{t})$ increases in the order of $\Theta(t)$ could be satisfied by gradient-descent based algorithm. Take Algorithm 1 as an example. For simplicity, let’s assume $a^{*}$ is far away from the boundary of $\mathcal{A}$ and projection never happens. It implies that $a^{\prime}_{t}=a_{t}+\delta u_{t},\forall t\in[T]$. This assumption is without loss of generality, since most of the proposed action $a^{\prime}_{t}$ will belong to the interior of the action space $\mathcal{A}$. This is because we start with $a_{0}=0$ and has small exploration size $\delta\sim\Theta(T^{-\alpha})$. Roughly speaking, we can express $\bar{V}_{t}=\lambda\mathbb{I}_{d}+\delta^{2}\sum^{t}_{s=1}u_{s}u^{\top}_{s}$, where $u_{s}$ is uniformly sampled from $\mathbb{S}^{d}$. Then applying Lemma A.2, we have $\lambda_{\min}(\bar{V}_{t})\sim\Theta(t)$ with high probability.
Lemma A.1 (Generalized Elliptical Lemma in Wang et al. (2023)).
Suppose $V_{0}\in\mathbb{R}^{d\times d}$ is any positive definite matrix, $a_{1},\ldots,a_{T}\in\mathbb{R}^{d}$ is a sequence of vectors with bounded $l_{2}$ norm and $\bar{V}_{t}:=V_{0}+\sum^{t}_{s=1}a_{s}a^{\top}_{s}$. Then for any $\rho\in[0,1]$, the following inequality holds with probability at least $1-\delta$
$\sum^{T}_{t=1}\left(\min\left\{1,\|a_{t}\|^{2}_{\bar{V}^{-1}_{t}}\right\}\right)^{1-\rho}\leq 2^{1-\rho}T^{\rho}\log^{1-\rho}\left(\frac{\det(\bar{V}_{t})}{\det(V_{0})}\right).$
Lemma A.2 (Proposition 1 in Li et al. (2017)).
Define $V_{t}:=\sum^{t}_{s=1}u_{s}u^{\top}_{s}$, where $u_{s}$ is drawn iid from some distribution $\nu$ with support on the unit ball, $\mathbb{B}^{d}$. Furthermore, let $\Sigma:=\mathbb{E}[u_{s}u^{\top}_{s}]$ be the second moment matrix. $B$ and $\delta>0$ are two positive constants. Then, there exist positive, universal constants $C_{1}$ and $C_{2}$ such that $\lambda_{\min}(V_{s})\geq B$ with probability at least $1-\delta$, as long as
$t\geq\left(\frac{C_{1}\sqrt{d}+C_{2}\sqrt{\log(1/\delta)}}{\lambda_{\min}(\Sigma)}\right)^{2}+\frac{2B}{\lambda_{\min}(\Sigma)}.$
B Missing Proofs in Theorem 1 and Proposition 1
B.1 Proof for Theorem 1: A Tight Lower Bound
@@
There exists an LIHF instance with a $\rho$-imperfect user (Def. 1) and strongly concave user utility such that any learner has to suffer $\text{Reg}_{T}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$ even with the knowledge of $\rho$. Moreover, this regret lower bound is tight, up to logarithmic terms. That is, for the same setting, there exists an online learning algorithm with $\text{Reg}_{T}\leq\tilde{O}(d\max\{\sqrt{T},T^{\rho}\})$.
Proof (Proof for Regret Lower Bound in Theorem 1). Before showing the regret lower bound proof for Theorem 1, let’s first establish the following Lemma B.1, which builds the connection between regret occurred under bandit reward feedback and dueling feedback. This proof is inspired by Theorem 6 in Yao et al. (2022a). To prove Lemma B.1, in addition to $\text{Reg}_{T}$, we introduce a new metric to measure the performance of algorithm under dueling feedback, which we coin functional regret $\text{Reg}^{\text{FO}}_{T}$, defined as follows.
@@
Lemma B.1.
Given a utility function $\mu$, if the regret lower bound for algorithm with bandit reward feedback is $\overline{\text{Reg}}$, then any learner $\mathcal{L}$ with dueling feedback has to occur regret $\text{Reg}^{\text{FO}}_{T}\geq 2\overline{\text{Reg}}$.
Proof  (Proof for Lemma B.1). We prove our claim by contradiction. Let $(a_{0,t},a_{1,t})$ be the pair of recommendation at round $t$. Suppose $\text{Reg}^{\text{FO}}_{T}<2\overline{\text{Reg}}$. As a result, at least one of the following inequality must hold:
$\begin{split}\mathbb{E}\left\{\sum^{T}_{t=1}f(a_{0,t})-f(a^{*})\right\}&\leq\overline{\text{Reg}};\\ \mathbb{E}\left\{\sum^{T}_{t=1}f(a_{1,t})-f(a^{*})\right\}&\leq\overline{\text{Reg}}.\end{split}$ (6)
Consider a principal who can observe the interaction between a user and the learner $\mathcal{L}$, then Then we can construct two algorithms $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$ as follows.
@@
2       Call algorithm $\mathcal{L}$ to generate two candidates $(a_{0,t},a_{1,t})$.
@@
4       Return the feedback $a_{*,t}$ to algorithm $\mathcal{L}$ and update $\mathcal{L}$.
@@
From Equation (6), we know that at least one of $\{\mathcal{L}_{0},\mathcal{L}_{1}\}$ achieves an expected regret lower than $\overline{\text{Reg}}$. However, we know all algorithm with bandit feedback with utility function $\mu$ has to occur regret at least $\overline{\text{Reg}}$, contradiction. Therefore, Lemma B.1 must hold, which completes the proof.
At the next step, we will prove Theorem 1 by discussing the value of $\rho$ case by case. Consider the scenario when $\rho\leq 0.5$. This is equivalent to say the total corruption level $C\leq\sqrt{T}$. In this scenario, since the utility function is strongly concave, applying Lemma B.2, we know $\overline{\text{Reg}}=0.02\min\left\{T,d\sqrt{T}\right\}$. By Lemma B.1, we get $\text{Reg}^{\text{FO}}_{T}\geq 0.04\min\left\{T,d\sqrt{T}\right\}$. Consider a linear link function $\sigma(x)=\frac{1}{2}+\frac{1}{2}x$, we have $\text{Reg}_{T}\geq 0.02\min\left\{T,d\sqrt{T}\right\}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, which completes the proof.
Lemma B.2 (Theorem 6 in (Shamir, 2013)).
Let the number of rounds $T$ be fixed. The for any learner $\mathcal{L}$, there exists a quadratic function of the form $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}$ which is minimized at $\theta$ and $\|\theta\|_{2}\leq 0.5$ such that
@@
Now, let’s focus on the scenario when $\rho>0.5$, which is equivalent to $C>\sqrt{T}$. In essence, we want to extend Lemma B.2 to the scenario in presence of adversarial corruption induced by $\rho$-imperfect user. If we can show the regret lower bound can be generalized to $\Omega(dT^{\rho})$, then applying the same technique which uses Lemma B.1 to connect regret suffered under bandit reward feedback and regret suffered under dueling feedback and choose linear link function to connect $\text{Reg}^{\text{FO}}_{T}$ and $\text{Reg}_{T}$, we will get the desired regret lower bound. The extension of Lemma B.2 is proven at section B.1.1.
B.1.1 Proof for Lemma 1: Regret Lower Bound for Strongly Concave Utility
@@
Assume that the action space $\mathcal{A}$ is contained in a $d$-dimensional unit ball. Consider the utility function $\mu$ in the form $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, where $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$ is a random vector. Under corruption induced by $\rho$-imperfect user (Def. 1), for any fixed number of rounds $T$ and $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, there exists a preference parameter $\theta$ such that for any learner $\mathcal{L}$ under reward feedback, even with the knowledge of $\rho$, has to suffer regret $\text{Reg}_{T}:=\mathbb{E}\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\}\geq\frac{d}{4}C_{\kappa}T^{\rho}.$
Proof  (Proof for Lemma 1). To prove Lemma 1, we aim to prove that there exists a corruption strategy such that the regret incurred by any learner $\mathcal{L}$ with corrupted dueling feedback is not less than $\Omega(dT^{\rho})$. The formulation of such a corruption strategy hinges on the observation that when the utility function $\mu_{\theta}$ adopts the quadratic form parameterized by $\theta$, specifically $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, which is both smooth and strongly concave, then the regret incurred by $\mathcal{L}$ is bounded below by the sum of the Kullback-Leibler (KL) divergence between the distribution of the corrupted reward feedback obtained at round $t$, denoted $\hat{v}_{t},t\in[1,T]$, conditioned on different possible values of $\theta$. (see Lemma B.4). Assume that $\theta$ is uniformly drawn from $\{-\beta,\beta\}^{d}$, given an action $a_{t}$, conditioned on $\theta_{i}>0$, the corrupted reward feedback $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}>0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)+\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}>0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{1}}}+\xi_{a_{t}}.$
@@
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}<0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)-\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}<0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{2}}}+\xi_{a_{t}}.$
We use $c_{t}(a_{t}|\theta_{i}>0)$ and $c_{t}(a_{t}|\theta_{i}<0)$ to empathize the fact that the magnitude and sign of corruption can be dependent on $\theta$. $\xi_{a_{t}}$ follows a standard Gaussian distribution. Therefore, $\hat{v}_{t}$ in Equation 2 follows $\text{N}(\mu_{1},1)$. $\hat{v}_{t}$ in Equation 3 follows $\text{N}(\mu_{2},1)$. By using Lemma B.3, we have
$\mathbb{D}_{t,i}=\text{D}_{\text{KL}}(\text{N}(\mu_{1},1)||\text{N}(\mu_{2},1))=(\mu_{1}-\mu_{2})^{2}.$
Lemma B.3 (KL Divergence for Normal Distribution).
@@
To optimize the lower bound in Lemma B.4, the adversary selects $c_{t}(a_{t}|\theta_{i}>0)$ and $c_{t}(a_{t}|\theta_{i}<0)$ to minimize the difference between $\mu_{1}$ and $\mu_{2}$. Consider the following corruption strategy: when $\theta_{i}>0$, set $c_{t}(a_{t}|\theta_{i}>0)=-\beta a_{t,i}$; when $\theta_{i}<0$, set $c_{t}(a_{t}|\theta_{i}<0)=\beta a_{t,i}$. This strategy ensures that $\mu_{1}=\mu_{2}$.
@@
Since $|a_{t,i}-\theta_{i}|\geq||a_{t,i}|-\beta|$ for all $i$, we obtain the second inequality. By utilizing the corruption constraint (7), we derive the last inequality, which is minimized when $\|a_{t}\|_{\infty}=\frac{C_{\kappa}T^{\rho-1}}{\beta}$ for all $t$. Together with Lemma B.4, we have:
@@
We select $\beta$ to optimize the lower bound in Equation (8). Since $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, we can set $\beta=\sqrt{C_{\kappa}}T^{\frac{\rho}{2}-\frac{1}{2}}$ to satisfy $\|\theta\|_{2}\leq 1$. Then we obtain $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq dC_{\kappa}T^{\rho}/4$. Notice that the corruption level each round is bounded by $\beta\|a_{t}\|_{\infty}$. Consider the scenario when the radius of the action space $\mathcal{A}$ is upper bounded by $\beta$, then the corruption budget each round is bounded by $C_{\kappa}t^{-1+\rho}$, which satisfies the corruption induced by $\rho$-imperfect user, which completes the proof.
Lemma B.4.
Let’s consider the utility function $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$. Let $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{T}$ be a sequence of corrupted reward feedback obtained by a learner $\mathcal{L}$. Then there exists a $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$, uniformly drawn from $\{-\beta,\beta\}^{d}$, such that the regret occurred by $\mathcal{L}$ is
@@
Proof  (Proof for Lemma B.4). Using the similar argument in Shamir (2013), assume that the leaner $\mathcal{L}$ is deterministic: $a_{t}$ is a deterministic function of the realized corrupted reward feedback $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{t-1}$ at $a_{1},a_{2},\ldots,a_{t-1}$. This assumption is without loss of generality, since any random learners can be seen as a randomization over deterministic learning algorithms. Thus a lower bound which holds uniformly for any deterministic $\mathcal{L}$ would also hold over a randomization. To lower bound Equation (9), we use Lemma B.5, which relates this to the question of how informative are the query values (as measured by Kullback-Leibler divergence) for determining the sign of $\theta$’s coordinates. Intuitively, the more similar the query values are, the smaller is the KL divergence and the harder it is to distinguish the true sign of $\theta_{i}$, leading to a larger lower bound. In addition, we are facing a powerful adversary who has the complete knowledge of the problem and is able to add corruption on the query value to make they are even more similar, which resulting a even smaller KL divergence, consequently, an even larger lower bound. Let $\bar{a}_{T}=\frac{1}{T}\sum^{T}_{t=1}a_{t}$ represent the average action, we have
@@
We get the second inequality by using the fact that $\mu_{\theta}$ is 1-strongly concave. We get the last inequality by using Lemma B.5, which completes the proof.
Lemma B.5 (Lemma 4 in (Shamir, 2013)).
Let $\theta$ be a random vector, none of those coordinates is supported on 00. Let $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{T}$ be a sequence of values obtained by a deterministic learner $\mathcal{L}$ returning a point $\bar{a}_{T}$ (so that the action $a_{t}$ is a deterministic function of $\hat{v}_{1},\ldots,\hat{v}_{t-1}$ and $\bar{a}_{T}$ is a deterministic function of $\hat{v}_{1},\ldots,\hat{v}_{T}$). Then we have
@@
We defer reader to Appendix C.3 for the proof of matching regret upper bound and the detailed description of the algorithm which achieves that upper bound.
B.2 Proof for Proposition 1: Regret Lower Bound for Linear Utility

### v2

@@ 6 Conclusion
In conclusion, this paper enhances the understanding of continuous dueling bandits by addressing $\rho$-Imperfect Human Feedback with concave utilities. We established fundamental regret lower bounds, introduced the RoSMID algorithm with nearly optimal performance, and developed a novel regret analysis framework that highlights the inherent efficiency-robustness trade-off in gradient-based algorithms. Our experimental results corroborate the theoretical predictions. A promising future direction is the development of an algorithm that achieves $\tilde{O}(d\sqrt{T}+dT^{\rho})$ regret for continuous dueling bandits with strongly concave utilities under arbitrary adversarial corruption. Additionally, incorporating contextual information to better model real-world scenarios could further enhance the applicability and robustness of these algorithms. We left these directions for future research.
A Proofs for Theorem 1
@@
There exists a $\rho$-Imperfect Human Feedback (see Definition 1), strongly concave utility function $\mu$, and link function $\sigma$ such that any learner has to suffer $\text{Reg}_{T}\geq\Omega\left(d\max\{\sqrt{T},T^{\rho}\}\right)$, even with the knowledge of $\rho$.
Proof  Before showing the regret lower bound proof for Theorem 1, let’s first establish the following Lemma 7, which builds the connection between regret incurred under bandit reward feedback and dueling feedback. This proof is inspired by Theorem 6 in Yao et al. (2022). To prove Lemma 7, in addition to $\text{Reg}_{T}$, we introduce a new metric to measure the performance of algorithm under dueling feedback, which we coin functional regret $\text{Reg}^{\text{FO}}_{T}$, defined as follows.
@@
Lemma 7.
Given a utility function $\mu$, if the regret lower bound for algorithm with bandit reward feedback is $\overline{\text{Reg}}$, then any learner with dueling feedback has to suffer regret $\text{Reg}^{\text{FO}}_{T}\geq 2\overline{\text{Reg}}$.
Proof  We prove our claim by contradiction. Let $(a_{0,t},a_{1,t})$ be the pair of recommendation at round $t$. Suppose $\text{Reg}^{\text{FO}}_{T}<2\overline{\text{Reg}}$. As a result, at least one of the following inequality must hold:
$\begin{split}\mathbb{E}\left\{\sum^{T}_{t=1}\mu(a^{*})-\mu(a_{0,t})\right\}&\leq\overline{\text{Reg}};\\ \mathbb{E}\left\{\sum^{T}_{t=1}\mu(a^{*})-\mu(a_{1,t})\right\}&\leq\overline{\text{Reg}}.\end{split}$ (4)
Consider a principal who can observe the interaction between a user and the learner $\mathcal{L}$, then we can construct two algorithms $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$ as follows.
@@
2       Call the learner $\mathcal{L}$ to generate two candidates $(a_{0,t},a_{1,t})$.
@@
4       Return user feedback to the learner $\mathcal{L}$ and update $\mathcal{L}$.
@@
From (4), we know that at least one of $\{\mathcal{L}_{0},\mathcal{L}_{1}\}$ achieves an expected regret lower than $\overline{\text{Reg}}$. However, we know all algorithm with bandit feedback with utility function $\mu$ has to occur regret at least $\overline{\text{Reg}}$, contradiction. Therefore, Lemma 7 must hold, which completes the proof.
Next, we prove Theorem 1. First, consider the case $\rho\leq 0.5$. Given that the utility function is strongly concave, by applying Lemma 8, we have $\overline{\text{Reg}}=0.02\min\{T,d\sqrt{T}\}$. From Lemma 7, we also obtain $\text{Reg}^{\text{FO}}_{T}\geq 0.04\min\{T,d\sqrt{T}\}$. Using the linear link function $\sigma(x)=\frac{1}{2}+\frac{1}{2}x$, it follows that $\text{Reg}_{T}\geq 0.02\min\{T,d\sqrt{T}\}\geq\Omega(d\max\{\sqrt{T},T^{\rho}\})$, completing the proof.
Lemma 8 (Theorem 6 in (Shamir, 2013)).
Let the number of rounds $T$ be fixed. Then for any learner, there exists a quadratic function of the form $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}$ which is minimized at $\theta$ and $\|\theta\|_{2}\leq 0.5$ such that
@@
Now, let’s focus on the scenario when $\rho>0.5$. In essence, we want to extend Lemma 8 to the scenario in presence of adversarial corruption induced by $\rho$-Imperfect Human Feedback. If we can show the regret lower bound can be generalized to $\Omega(dT^{\rho})$, then applying the same technique which uses Lemma 7 to connect regret suffered under bandit reward feedback and regret suffered under dueling feedback and choose linear link function to connect $\text{Reg}^{\text{FO}}_{T}$ and $\text{Reg}_{T}$, we will get the desired regret lower bound. The extension of Lemma 8 is proven at section A.1.
A.1 Proof for Lemma 1
@@
Consider bandit learning with direct reward feedback, where reward $r(a):=\mu(a)+\epsilon$, $\epsilon$ follows standard normal distribution. The action space $\mathcal{A}$ is contained in a $d$-dimensional unit ball, and $\mu$ is $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, with $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$ Under $\rho$-Imperfect Human Feedback (Definition 1), for any $T$ and $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, there exists a $\theta$ such that for any learner under direct reward feedback, even with the knowledge of $\rho$, has to suffer regret $\text{Reg}_{T}:=\mathbb{E}\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\}\geq\frac{d}{4}C_{\kappa}T^{\rho}.$
Proof  To prove Lemma 1, we aim to prove that there exists a corruption strategy such that the regret incurred by any learner with imperfect dueling feedback is not less than $\Omega(dT^{\rho})$. The formulation of such a corruption strategy hinges on the observation that when the utility function $\mu_{\theta}$ adopts the quadratic form parameterized by $\theta$, specifically $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, which is both smooth and strongly concave, then the regret incurred by the learner is bounded below by the sum of the Kullback-Leibler (KL) divergence between the distribution of the corrupted reward feedback obtained at round $t$, denoted $\hat{v}_{t},t\in[T]$, conditioned on different possible values of $\theta$. (see Lemma 10). Assume that $\theta$ is uniformly drawn from $\{-\beta,\beta\}^{d}$, given an action $a_{t}$, conditioned on $\theta_{i}>0$, the corrupted reward feedback $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}>0)+\xi=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)+\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}>0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{1}}}+\xi.$ (5)
@@
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}<0)+\xi=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)-\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}<0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{2}}}+\xi.$ (6)
We use $c_{t}(a_{t}|\theta_{i}>0)$ and $c_{t}(a_{t}|\theta_{i}<0)$ to empathize the fact that the magnitude and sign of corruption can be dependent on $\theta$. $\xi$ follows a standard Gaussian distribution. Therefore, $\hat{v}_{t}$ in Equation 5 follows $\text{N}(\mu_{1},1)$. $\hat{v}_{t}$ in Equation 6 follows $\text{N}(\mu_{2},1)$. By using Lemma 9, we have
$\text{D}_{\text{KL}}(\text{N}(\mu_{1},1)||\text{N}(\mu_{2},1))=(\mu_{1}-\mu_{2})^{2}.$
Lemma 9 (KL Divergence for Normal Distribution).
@@
To optimize the lower bound in Lemma 10, the adversary selects $c_{t}(a_{t}|\theta_{i}>0)$ and $c_{t}(a_{t}|\theta_{i}<0)$ to minimize the difference between $\mu_{1}$ and $\mu_{2}$. Consider the following corruption strategy: when $\theta_{i}>0$, set $c_{t}(a_{t}|\theta_{i}>0)=-\beta a_{t,i}$; when $\theta_{i}<0$, set $c_{t}(a_{t}|\theta_{i}<0)=\beta a_{t,i}$. This strategy ensures that $\mu_{1}=\mu_{2}$.
@@
Since $|a_{t,i}-\theta_{i}|\geq||a_{t,i}|-\beta|$ for all $i$, we obtain the second inequality. By utilizing the corruption constraint (7), we derive the last inequality, which is minimized when $\|a_{t}\|_{\infty}=\frac{C_{\kappa}T^{\rho-1}}{\beta}$ for all $t$. Together with Lemma 10, we have:
@@
We select $\beta$ to optimize the lower bound in Equation (8). Since $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, we can set $\beta=\sqrt{C_{\kappa}}T^{\frac{\rho}{2}-\frac{1}{2}}$ to satisfy $\|\theta\|_{2}\leq 1$. Then we obtain $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq dC_{\kappa}T^{\rho}/4$. Notice that the corruption level each round is bounded by $\beta\|a_{t}\|_{\infty}$. Consider the scenario when the radius of the action space $\mathcal{A}$ is upper bounded by $\beta$, then the corruption budget each round is bounded by $C_{\kappa}t^{-1+\rho}$, which satisfies the definition of $\rho$-Imperfect Human Feedback, which completes the proof.
Lemma 10.
Let’s consider the utility function $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$. Let $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{T}$ be a sequence of corrupted reward feedback obtained by a learner. Then there exists a $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$, uniformly drawn from $\{-\beta,\beta\}^{d}$, such that the regret suffered by the learner is
@@
Proof  Using the similar argument in Shamir (2013), assume that the learner is deterministic: $a_{t}$ is a deterministic function of the realized corrupted reward feedback $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{t-1}$ at $a_{1},a_{2},\ldots,a_{t-1}$. This assumption is without loss of generality, since any random learners can be seen as a randomization over deterministic learning algorithms. Thus a lower bound which holds uniformly for any deterministic learner would also hold over a randomization. To lower bound (9), we use Lemma 11, which relates this to the question of how informative are the query values (as measured by Kullback-Leibler divergence) for determining the sign of $\theta$’s coordinates. Intuitively, the more similar the query values are, the smaller is the KL divergence and the harder it is to distinguish the true sign of $\theta_{i}$, leading to a larger lower bound. In addition, we are facing a powerful adversary who has the complete knowledge of the problem and is able to add corruption on the query value to make they are even more similar, which resulting a even smaller KL divergence, consequently, an even larger lower bound. Let $\bar{a}_{T}:=\frac{1}{T}\sum^{T}_{t=1}a_{t}$ represent the average action, we have
@@
We get the second inequality by using the fact that $\mu_{\theta}$ is 1-strongly concave. We get the last inequality by using Lemma 11, which completes the proof.
Lemma 11 (Lemma 4 in (Shamir, 2013)).
Let $\theta$ be a random vector, none of those coordinates is supported on 00. Let $\hat{v}_{1},\hat{v}_{2},\ldots,\hat{v}_{T}$ be a sequence of values obtained by a deterministic learner returning a point $\bar{a}_{T}$ (so that the action $a_{t}$ is a deterministic function of $\hat{v}_{1},\ldots,\hat{v}_{t-1}$ and $\bar{a}_{T}$ is a deterministic function of $\hat{v}_{1},\ldots,\hat{v}_{T}$). Then we have
@@
A.2 Proof for Proposition 1: